The Evergreen hotel, a 200-room lodging facility, uses regression analysis to forecast dining rooms meals and revenues. You, as a GM, has set the regression equations as follows:

Breakfast: y = 50 + 0.42x

Lunch: y = 200 + 0.21x

where, “y” equals forecasted meals and “x” is the #of hotel guests. Further, daily average guest check in the hotel’s dining room is expected to be as follows: Breakfast: $3.25 Lunch: $6.50 Dinner: $12.95 Based on the information given, forecast the daily revenue for each meal period when the occu-pancy is 85.00% and when the average occupancy per room is expected to be 1.58. Also, what is the forecasted daily total dining room revenue of the Evergreen hotel (assume that all of the hotel guests eat breakfast, lunch, and dinner at the hotel)?
Answer #1
Total room        200.00
Occu 85%
Total occupied room        170.00
Guests per room             1.58
Total guests        268.60
Breakfast y= 50+ (0.42 x 268.6)
          163
Lunch y= 200+ (0.21 x 268.6)
          256
Revenue
Breakfast per plate          3.25
Total breakfast revenue = 163 X 3.25
          530 Answer for breakfast revenue per day
Lunch per plate          6.50
Total lunch revenue = 256 X 6.50
       1,664 Answer for lunch revenue per day
If all guests take all 3 meal from the hotel, then the revenue would be -
Total breakfast revenue = 268.6 X 3.25 Total guests
          873
Lunch per plate          6.50 Total guests
Total lunch revenue = 268.6 X 6.50
       1,746
Dinner per plate       12.95 Total guests
Total dinner revenue = 268.6 X 12.95
       3,478
Total revenue =873+1746+3478
       6,097 Answer for the last question